A chemist needs to make 30 ounces of a 25% alcohol solution by mixing together a 15% alcohol solution with a 40% alcohol solution. How much of each should they use?
Answer by Maths68(1474)

Solution A
Amount = x
Concentration =15% = 0.15
Solution B
Amount = 30-x
Concentration =40% = 0.40
Resultant Solution
Amount = 30 ounce
Concentration =25% = 0.25

[Amount Solution A * Concentration A] + [Amount Solution B * Concentration of B]+ [Amount Solution B * Concentration of B] = Amount of Resultant * Concentration of resultant
x(0.15)+(30-x)(0.40)=(30)(0.25)
0.15x+12-0.4x=7.5
-0.25x=7.5-12
-0.25x=-4.5
-0.25x/-0.25=-4.5/-0.25
x=18




Solution A
Amount = x = 18
Concentration =15% = 0.15
Solution B
Amount = 30-x = 30-18 = 12
Concentration =40% = 0.40

18 ounces of 15% alcohol solution with a 12 ounces of 40% alcohol solution will be mixed to get 30 ounces of a 25% alcohol solution.